# On the kernels of the pro-$p$ outer Galois representations associated to once-punctured CM elliptic curves

### v2

@@ 4.1. Construction of elements
in the two-variable graded quotient $\mathrm{gr}^{\boldsymbol{m}+(1,0)}_{1}G$. However, by Corollary 3.15, we also have
@@
in the same graded quotient. Since the element $\chi_{1}(\delta_{1})\in\mathbb{Z}_{p}^{\times}$ is of order $p-1$, we have $\sigma_{\boldsymbol{m}+(p-1,0)}\in F^{\boldsymbol{m}+(2,0)}G$. By applying the same argument to graded quotients
$\mathrm{gr}^{\boldsymbol{m}+(2,0)}_{1}G,\dots,\mathrm{gr}^{\boldsymbol{m}+(p-2,0)}_{1}G,$
@@
For every $m\geq\lvert\boldsymbol{m}_{0}\rvert$ such that $m\equiv\lvert\boldsymbol{m}_{0}\rvert\bmod p-1$, let $A^{\boldsymbol{m}_{0},m}$ be the image of $F^{m}G$ in $A^{\boldsymbol{m}_{0}}$. Note the image of $F^{m+1}G$ in $A^{\boldsymbol{m}_{0},m}$ coincides with $A^{\boldsymbol{m}_{0},m+(p-1)}$. In fact, assume that there is an integer $11\leq i<p-$ 1 such that the image of $F^{m+i}G$ properly contains $A^{\boldsymbol{m}_{0},m+(p-1)}$, and take $i$ to be maximal among such integers. Then we would have a nontrivial homomorphism
@@
In the following, we prove the first half of the claim by induction on $m$. We identify the ring $\Lambda=\mathbb{Z}_{p}[[\mathrm{Gal}(K(p^{\infty})/K(p))]]$ with the power series ring $\mathbb{Z}_{p}[[S_{1},S_{2}]]$ via $S_{i}=\gamma_{i}-1$ for $i=1,2$.
First, we know that the claim holds for $m=|\boldsymbol{m}_{0}|$. We prove the claim for $m+(p-1)$, assuming it is true for $m$. Take an arbitrary element $x\in A^{\boldsymbol{m}_{0},m+(p-1)}$. By induction hypothesis, the element $x$ can be written as
@@
Note that $\kappa_{\boldsymbol{n}}(g_{\boldsymbol{n}})$ generates $\kappa_{\boldsymbol{n}}(F^{\lvert\boldsymbol{n}\rvert}G)$ by induction hypothesis. Moreover, the submodule $\kappa_{\boldsymbol{n}}(F^{\lvert\boldsymbol{n}\rvert}G)\subset\mathbb{Z}_{p}(\boldsymbol{n})$ is nonzero by Theorem 2.6 (2) and Corollary A.3. Therefore we have $\kappa_{\boldsymbol{n}}(g_{\boldsymbol{n}})\neq 0$, implying that $f_{\boldsymbol{n}}(\chi_{1}^{n_{1}}(\gamma_{1})-1,\chi_{2}^{n_{2}}(\gamma_{2})-1)=0$. Equivalently, the power series $f_{\boldsymbol{n}}(S_{1},S_{2})$ is contained in the ideal $(S_{1}-\chi_{1}^{n_{1}}(\gamma_{1})+1,S_{2}-\chi_{2}^{n_{2}}(\gamma_{2})+1)$. Since we have
@@
4.2. Group-theoretic lemmas
@@
for some $a_{i,j}\in\mathbb{Z}_{p}$ for every $j\geq 1$. Denote the normal closure of $\{x_{i}\}_{i\geq 1}$ in $\mathcal{F}$ by $H$. Then the following two assertions hold:
@@
is pro-$p$, since the kernel of each $\mathrm{Aut}(\mathcal{K}/\mathcal{K}_{i,j})\to\mathrm{Aut}(M/M_{i,j})$ is so. Therefore, to prove the claim, it suffices to show that the homomorphism ^$\hat{\mathbb{Z}}\to\mathrm{Aut}_{\rm{fil}}(M)$ corresponding to the image of $\phi$ factors through $\mathbb{Z}_{p}$.
@@
$\phi^{p^{n}}\in\ker\left(\mathrm{Aut}_{\rm{fil}}(M)\to\mathrm{Aut}(M/M_{p^{n},p^{n}})\right)$
for every $n\geq 1$. In fact, a direct computation shows that $\phi^{p}(\tilde{x}_{i,j})=\tilde{x}_{i,j+p}+\tilde{x}_{i,j}$ in $M$ for every $i,j$, and iterating $\phi^{p}$ verifies the claim. Hence the image of $\phi$ in $\mathrm{Aut}_{\rm{fil}}(M)=\mathrm{Aut}(M/M_{p^{n},p^{n}})$ is a pro-$p$ group. It follows that the map $\mathbb{Z}\to\mathrm{Aut}_{\rm{fil}}(\mathcal{K})$ corresponding to $\phi$ naturally extends to a continuous homomorphism $\mathbb{Z}_{p}\to\mathrm{Aut}_{\rm{fil}}(\mathcal{K})$, and we can take the associated semi-direct product $\mathcal{K}\rtimes\mathbb{Z}_{p}$.
@@
Let $r\geq 1$ be an integer and $\mathcal{F}$ be a pro-$p$ group generated by $y_{1},y_{2}$ and $\{x_{i}\}_{1\leq i\leq r}$. For each $1\leq i\leq r$, let $x_{i,(0,0)}\coloneqq x_{i}$ and we inductively define
@@
for some $\beta_{i,j}\in\mathbb{Z}_{p}$ and each $j\geq 0$. Let $H$ denote the normal closure of $\{x_{i}\}_{1\leq i\leq r}$ and $z_{(0,0)}$. Then $H$ is strongly generated by $\{x_{i,(j,k)}\}_{\begin{subarray}{c}1\leq i\leq r,\\ j,k\geq 0\end{subarray}}$ and $\{z_{i,j}\}_{i,j\geq 0}$.
@@
Since $[\tilde{F}^{m}\mathcal{G},\tilde{F}^{1}\mathcal{G}]\subset\tilde{F}^{m}\mathcal{G}$, it follows that $\tilde{F}^{m+1}\mathcal{G}\subset\tilde{F}^{m}\mathcal{G}$ for every $m\geq 1$. Apparently, the filtration $\{\tilde{F}^{m}\mathcal{G}\}_{m\geq 1}$ defines a descending central filtration on $\mathcal{G}$ satisfying the condition ($\mathrm{i}$) in the assertion of (1). Moreover, if $\{F^{m}\mathcal{G}\}_{m\geq 1}$ is a descending central filtration satisfying ($\mathrm{i}$), then we have $F^{1}\mathcal{G}=\tilde{F}^{1}\mathcal{G}=\mathcal{G}$. By induction on $m$, it follows that $\tilde{F}^{m}\mathcal{G}\subset F^{m}\mathcal{G}$ holds for every $m\geq 1$. Hence the filtration $\{\tilde{F}^{m}\mathcal{G}\}_{m\geq 1}$ also satisfies ($\mathrm{ii}$). The uniqueness is clear. The proof of the assertion of (2) is similar to that of [Iha02, p.263, 5].
@@
We claim that the image of $\tilde{F}^{m}\mathcal{G}$ in $F_{n}$ contained in the $r_{m}$-th component $F_{n}(r_{m})$ of the descending central series of $F_{n}$, where an integer $r_{m}$ is defined by
@@
for every $m\geq 1$. Once this claim is proved, then the image of $\cap_{m\geq 1}\tilde{F}^{m}\mathcal{G}$ in $F_{n}$ is contained in the intersection $\cap_{m\geq n}F_{n}(r_{m})=\{1\}$, and hence the assertion follows.
The claim holds for every $m\leq 2n-1$. Assume that the claim also holds for every $m\leq kn-1$ for some $k\geq 2$ and we prove the claim for $m=kn,kn+1,\dots(k+1)n-1$ in order. Write $m=kn+r$ for some $10\leq r\leq n-$ 1.
Let $m^{\prime},m^{\prime\prime}$ be positive integers less than $m$ satisfying $m^{\prime}+m^{\prime\prime}\geq m$, and write $m^{\prime}=k^{\prime}n+r^{\prime}$, $m^{\prime}=k^{\prime\prime}n+r^{\prime\prime}$ for some $0\leq k^{\prime},k^{\prime\prime}\leq k$ and $10\leq r^{\prime},r^{\prime\prime}\leq n-$ 1. Since
@@
Therefore, the image of $[\tilde{F}^{m^{\prime}}\mathcal{G},\tilde{F}^{m^{\prime\prime}}\mathcal{G}]$ is contained in
$[F_{n}(r_{m^{\prime}}),F_{n}(r_{m^{\prime\prime}})]\subset F_{n}(r_{m^{\prime}}+r_{m^{\prime\prime}})\subset F_{n}(r_{m}).$
@@
The next lemma gives an explicit set of generators of $G=\mathrm{Gal}(\Omega/K(p))$:
@@
We use Lemma 4.2 and definitions of $\sigma_{\boldsymbol{m}}$. It follows that the image of $\sigma_{\boldsymbol{m}}$ in $A=\mathrm{Gal}(\Omega/K(p^{\infty}))^{{\rm ab}}$ generates $A^{\boldsymbol{m}}$ for every $\boldsymbol{m}\in I_{0}\setminus\{(p,1),(1,p),(p-1,p-1)\}$. Moreover, the image of $\sigma_{(p,1)}$ and $\sigma_{(1,p)}$ also generates $A^{(1,1)}$. Finally, recall that $\sigma_{(p-1,p-1)}$ coincides with the commutator $[\gamma_{1},\gamma_{2}]$. We have proved in the proof of Lemma 4.4 that this element generates $A^{(p-1,p-1)}$.
Hence it follows that the $\Lambda$-module $A$ is generated by the image of $\{\sigma_{\boldsymbol{m}}\}_{\boldsymbol{m}\in I_{0}}$, which is equivalent to saying that $\mathrm{Gal}(\Omega/K(p^{\infty}))$ is normally generated by $\{\sigma_{\boldsymbol{m}}\}_{\boldsymbol{m}\in I_{0}}$ (as a normal subgroup of $G$). Hence $\gamma_{1}$, $\gamma_{2}$ and $\{\sigma_{\boldsymbol{m}}\}_{\boldsymbol{m}\in I_{0}\setminus\{(p-1,p-1)\}}$ generates the group $G$, as desired. ∎
@@
Now we recall the discussion occurring at constructing $\sigma_{\boldsymbol{m}}$ and $g_{\boldsymbol{m}}$ for $\boldsymbol{m}\in I$ such that $\boldsymbol{m}\equiv\boldsymbol{1}\bmod p-1$. There were two ways to construct them, starting from $\sigma_{(p,1)}$ and $\sigma_{(1,p)}$, but they yield the same elements on $A=\mathrm{Gal}(\Omega/K(p^{\infty}))^{{\rm ab}}$. Hence, by applying Corollary 4.10 to $\mathcal{F}=G$, $y_{1}=\gamma_{1}$, $y_{2}=\gamma_{2}$ and $\{x_{i}\}_{1\leq i\leq r}=\{g_{\boldsymbol{m}}\}_{\boldsymbol{m}\in I_{0}}$, it follows that $F^{1}G=\mathrm{Gal}(\Omega/K(p^{\infty}))$ is strongly generated by $\{g_{\boldsymbol{m}}\}_{\boldsymbol{m}\in I}$. By Lemma 4.7, it follows that $\{\sigma_{\boldsymbol{m}}\}_{\boldsymbol{m}\in I}$ also strongly generates $F^{1}G$.
We claim that the projection $F^{1}G\to F^{1}G^{\ast}$ is an isomorphism, which is equivalent to the desired equality $\Omega=\Omega^{\ast}$. Let $\mathcal{G}$ be a free pro-$p$ group on the set $\{\tilde{\sigma}_{\boldsymbol{m}}\}_{\boldsymbol{m}\in I}$, $\varphi\colon\mathcal{G}\to F^{1}G$ a surjective homomorphism sending $\tilde{\sigma}_{\boldsymbol{m}}$ to $\sigma_{\boldsymbol{m}}$ for every $\boldsymbol{m}\in I$ and $\varphi^{\ast}\colon\mathcal{G}\to F^{1}G^{\ast}$ the composite of $\varphi$ with the projection $F^{1}G\to F^{1}G^{\ast}$.
The group $\mathcal{G}$ comes equipped with a descending central filtration $\{F^{m}\mathcal{G}\}_{m\geq 1}$ induced by $\{F^{m}G^{\ast}\}_{m\geq 1}$ through the map $\varphi^{\ast}$. Note that the graded Lie algebra associated to $\{F^{m}\mathcal{G}\}_{m\geq 1}$ is naturally isomorphic to $\mathfrak{g}$, and we have
$\ker(\varphi^{\ast})=(\varphi^{\ast})^{-1}(\cap_{m\geq 1}F^{m}G^{\ast})=\cap_{m\geq 1}F^{m}\mathcal{G}.$
Hence, to show the injectivity of the map $\varphi^{\ast}$, it suffices to prove that the intersection of the induced filtration on $\mathcal{G}$ is trivial. By Lemma 4.11 (1), the group $F^{m}\mathcal{G}$ contains $\tilde{F}^{m}\mathcal{G}$ for every $m\geq 1$. Hence we have the following commutative diagram:
@@
Since we assume the analogue of the Deligne-Ihara conjecture (Conjecture 2.8), it holds that $\{\tilde{\sigma}_{\boldsymbol{m}}\}_{\boldsymbol{m}\in I}$ freely generates $\bigoplus_{m\geq 1}(F^{m}\mathcal{G}/F^{m+1}\mathcal{G})\otimes\mathbb{Q}_{p}$ as a graded Lie algebra. However, by Lemma 4.11 (2), the Lie algebra $\bigoplus_{m\geq 1}(\tilde{F}^{m}\mathcal{G}/\tilde{F}^{m+1}\mathcal{G})\otimes\mathbb{Q}_{p}$ is also generated by $\{\tilde{\sigma}_{\boldsymbol{m}}\}_{\boldsymbol{m}\in I}$, which forces the upper horizontal arrow in the diagram to be an isomorphism. As a consequence, it holds that the lower horizontal arrow in the diagram is injective. As is observed by induction on $m$, this injectivity is equivalent to saying that two filtrations $\{\tilde{F}^{m}\mathcal{G}\}_{m\geq 1}$ and $\{F^{m}\mathcal{G}\}_{m\geq 1}$ coincide with each other. Hence by Lemma 4.11 (3), we have
$\cap_{m\geq 1}F^{m}\mathcal{G}=\cap_{m\geq 1}\tilde{F}^{m}\mathcal{G}=\{1\},$
showing the injectivity of the map $\varphi^{\ast}\colon\mathcal{G}\to F^{1}G^{\ast}$. Hence the map $\varphi^{\ast}$ is an isomorphism, so is the natural surjection $F^{1}G\to F^{1}G^{\ast}$, as desired. ∎
Remark 4.13.
@@
(1) As a byproduct of the proof, the elements $\{\sigma_{\boldsymbol{m}}\}_{\boldsymbol{m}\in I}$ freely generate the graded Lie algebra $\mathfrak{g}$ over $\mathbb{Z}_{p}$, by Lemma 4.11 (2). In other words, under the assumption of Theorem 2.12, the integral version of Conjecture 2.8 holds. In the same way that the integral version of the Deligne-Ihara conjecture is related to generalized Greenberg’s conjecture for $\mathbb{Q}(\mu_{p})$ [Sha02, Theorem 1.3], the integral version of Conjecture 2.8 is related to generalized Greenberg’s conjecture for $K(p)$. We hope to discuss this topic in a future paper.
@@
(2) As another byproduct, the Galois group $\mathrm{Gal}(\Omega/K(p^{\infty}))$ is found to be a free pro-$p$ group generated by $\{\sigma_{\boldsymbol{m}}\}_{\boldsymbol{m}\in I}$. This consequence is consistent with Remark 4.3 (2).
Appendix A

### v3

@@ 4.1. Construction of elements
in the two-variable graded quotient $\operatorname{Gr}^{\boldsymbol{m}+(1,0)}_{1}G$. However, by Corollary 3.19, we also have
@@
in the same graded quotient. Since the element $\chi_{1}(\delta_{1})\in\mathbb{Z}_{p}^{\times}$ is of order p1p-1, we have $\sigma_{\boldsymbol{m}+(p-1,0)}\in F^{\boldsymbol{m}+(2,0)}G$. By applying the same argument to graded quotients
$\operatorname{Gr}^{\boldsymbol{m}+(2,0)}_{1}G,\dots,\operatorname{Gr}^{\boldsymbol{m}+(p-2,0)}_{1}G,$
@@
For every $m\geq\lvert\boldsymbol{m}_{0}\rvert$ such that $m\equiv\lvert\boldsymbol{m}_{0}\rvert\bmod p-1$, let $A^{\boldsymbol{m}_{0},m}$ be the image of $F^{m}G$ in $A^{\boldsymbol{m}_{0}}$. Note the image of $F^{m+1}G$ in $A^{\boldsymbol{m}_{0},m}$ coincides with $A^{\boldsymbol{m}_{0},m+(p-1)}$. In fact, assume that there is an integer $1\leq i<p-1$ such that the image of $F^{m+i}G$ properly contains $A^{\boldsymbol{m}_{0},m+(p-1)}$, and take $i$ to be maximal among such integers. Then we would have a nontrivial homomorphism
@@
In the following, we prove the first half of the claim by induction on $m$. We identify the ring $\Lambda=\mathbb{Z}_{p}[[\operatorname{Gal}(K(p^{\infty})/K(p))]]$ with the power series ring $\mathbb{Z}_{p}[[S_{1},S_{2}]]$ via $S_{i}=\gamma_{i}-1$ for $i=1,2$.
First, we know that the claim holds for $m=|\boldsymbol{m}_{0}|$. We prove the claim for m+(p1)m+(p-1), assuming it is true for $m$. Take an arbitrary element $x\in A^{\boldsymbol{m}_{0},m+(p-1)}$. By induction hypothesis, the element $x$ can be written as
@@
Note that $\kappa_{\boldsymbol{n}}(g_{\boldsymbol{n}})$ generates $\kappa_{\boldsymbol{n}}(F^{\lvert\boldsymbol{n}\rvert}G)$ by induction hypothesis. Moreover, the submodule $\kappa_{\boldsymbol{n}}(F^{\lvert\boldsymbol{n}\rvert}G)\subset\mathbb{Z}_{p}(\boldsymbol{n})$ is nonzero by Lemma 4.1. Therefore we have $\kappa_{\boldsymbol{n}}(g_{\boldsymbol{n}})\neq 0$, implying that $f_{\boldsymbol{n}}(\chi_{1}^{n_{1}}(\gamma_{1})-1,\chi_{2}^{n_{2}}(\gamma_{2})-1)=0$. Equivalently, the power series $f_{\boldsymbol{n}}(S_{1},S_{2})$ is contained in the ideal $(S_{1}-\chi_{1}^{n_{1}}(\gamma_{1})+1,S_{2}-\chi_{2}^{n_{2}}(\gamma_{2})+1)$. Since we have
@@
Remark 4.9.1.
In our construction of $\{\sigma_{\boldsymbol{m}}\}_{\boldsymbol{m}\in I}$, we assume the two assumptions
  1. (1)
    The class number of $K(p)$ is not divisible by $p$, and
  2. (2)
    There are exactly two primes of $K(p^{2})$ above $p$
to apply Lemma 4.3. However, it is clear from our construction that one can remove these two assumptions, if there exists a quotient $\bar{A}$ of the $\Lambda$-module $A$ that satisfies the following properties:
  1. (a)
    There exists a surjective homomorphism $\Lambda^{[K(p):K]-1}\oplus\operatorname{Ann}(\mathbb{Z}_{p}(1))\to\bar{A}$, and
  2. (b)
    The character $\kappa_{\boldsymbol{m}}$ factors through $\bar{A}$.
For example, let $\bar{A}$ be the Galois group of a Kummer extension of $K(p^{\infty})$ obtained by adjoining all $p$-powerth roots of elliptic units in $K(p^{\infty})$. Then it satisfies (b), but we do not know whether it also satisfies (a) or not.
4.2. Group-theoretic lemmas and end of proof of main result
@@
for some $a_{i,j}\in\mathbb{Z}_{p}$ for every $j\geq 1$. Denote the normal closure of $\{x_{i}\}_{i\geq 1}$ in $\mathcal{F}$ by $H$. The following assertions hold:
@@
is pro-$p$, since the kernel of each $\operatorname{Aut}(\mathcal{K}/\mathcal{K}_{i,j})\to\operatorname{Aut}(M/M_{i,j})$ is so. Therefore, to prove the claim, it suffices to show that the homomorphism $\hat{\mathbb{Z}}\to\operatorname{Aut}_{\rm{fil}}(M)$ corresponding to the image of $\phi$ factors through $\mathbb{Z}_{p}$.
@@
$\phi^{p^{n}}\in\operatorname{Ker}\left(\operatorname{Aut}_{\rm{fil}}(M)\to\operatorname{Aut}(M/M_{p^{n},p^{n}})\right)$
for every $n\geq 1$. In fact, a direct computation shows that $\phi^{p}(\tilde{x}_{i,j})=\tilde{x}_{i,j+p}+\tilde{x}_{i,j}$ in $M$ for every $i,j$, and iterating $\phi^{p}$ verifies the claim. Hence the image of $\phi$ in $\operatorname{Aut}_{\rm{fil}}(M)=\operatorname{Aut}(M/M_{p^{n},p^{n}})$ is a pro-$p$ group. It follows that the map $\mathbb{Z}\to\operatorname{Aut}_{\rm{fil}}(\mathcal{K})$ corresponding to $\phi$ naturally extends to a continuous homomorphism $\mathbb{Z}_{p}\to\operatorname{Aut}_{\rm{fil}}(\mathcal{K})$, and we can take the associated semi-direct product $\mathcal{K}\rtimes\mathbb{Z}_{p}$.
@@
Let $r$ be a positive integer and $\mathcal{F}$ a pro-$p$ group generated by $y_{1},y_{2}$ and $\{x_{i}\}_{1\leq i\leq r}$. For each $1\leq i\leq r$, let $x_{i,(0,0)}\coloneqq x_{i}$ and we inductively define
@@
for some $\beta_{i,j}\in\mathbb{Z}_{p}$ and each $j\geq 0$. Let $H$ denote the normal closure of $\{x_{i}\}_{1\leq i\leq r}$ and $z_{(0,0)}$ inside $\mathcal{F}$. Then $H$ is strongly generated by $\{x_{i,(j,k)}\}_{\begin{subarray}{c}1\leq i\leq r,\\ j,k\geq 0\end{subarray}}$ and $\{z_{i,j}\}_{i,j\geq 0}$.
@@
Since $\tilde{F}^{m}\mathcal{G},\tilde{F}^{1}\mathcal{G}]\subset\tilde{F}^{m}\mathcal{G}$, it follows that $\tilde{F}^{m+1}\mathcal{G}\subset\tilde{F}^{m}\mathcal{G}$ for every $m\geq 1$. Apparently, the filtration $\{\tilde{F}^{m}\mathcal{G}\}_{m\geq 1}$ defines a descending central filtration on $\mathcal{G}$ that satisfies the condition ($\mathrm{i}$) in the assertion of (1). Now let $\{F^{m}\mathcal{G}\}_{m\geq 1}$ be an arbitrary descending central filtration on $\mathcal{G}$ that satisfies ($\mathrm{i}$). We have $F^{1}\mathcal{G}=\tilde{F}^{1}\mathcal{G}=\mathcal{G}$. By induction on $m$, it immediately follows that $\tilde{F}^{m}\mathcal{G}\subset F^{m}\mathcal{G}$ holds for every $m\geq 1$. Hence the filtration $\{\tilde{F}^{m}\mathcal{G}\}_{m\geq 1}$ also satisfies ($\mathrm{ii}$). The uniqueness is clear. The proof of the assertion of (2) is similar to that of [Iha02, p.263, 5].
@@
We claim that the image of $\tilde{F}^{m}\mathcal{G}$ in $F_{n}$ contained in the $r_{m}$-th component $F_{n}(r_{m})$ of the descending central series of $F_{n}$, where $r_{m}$ is defined by
@@
for every $m\geq 1$. Once the claim is obtained, the image of $\cap_{m\geq 1}\tilde{F}^{m}\mathcal{G}$ in $F_{n}$ is contained in the intersection $\cap_{m\geq n}F_{n}(r_{m})=\{1\}$, and the assertion follows.
The claim trivially holds for every $m\leq 2n-1$. Assume that the claim also holds for every $m\leq kn-1$ for some $k\geq 2$. We prove the claim for $m=kn,kn+1,\dots(k+1)n-1$ in order. Write $m=kn+r$ for some $0\leq r\leq n-1$.
Let $m^{\prime},m^{\prime\prime}$ be positive integers which are less than $m$ and $m^{\prime}+m^{\prime\prime}\geq m$, and write $m^{\prime}=k^{\prime}n+r^{\prime}$, $m^{\prime}=k^{\prime\prime}n+r^{\prime\prime}$ for some $0\leq k^{\prime},k^{\prime\prime}\leq k$ and $0\leq r^{\prime},r^{\prime\prime}\leq n-1$. Since
@@
Therefore, the image of $\tilde{F}^{m^{\prime}}\mathcal{G},\tilde{F}^{m^{\prime\prime}}\mathcal{G}]$ is contained in
$F_{n}(r_{m^{\prime}}),F_{n}(r_{m^{\prime\prime}})]\subset F_{n}(r_{m^{\prime}}+r_{m^{\prime\prime}})\subset F_{n}(r_{m}).$
@@
The next lemma gives an explicit set of generators of $G=\operatorname{Gal}(\Omega/K(p))$:
@@
We use Lemma 4.3 and definitions of $\sigma_{\boldsymbol{m}}$. It follows that the image of $\sigma_{\boldsymbol{m}}$ in $A=\operatorname{Gal}(\Omega/K(p^{\infty}))^{{\rm ab}}$ generates the $\Lambda$-module $A^{\boldsymbol{m}}$ for every $\boldsymbol{m}\in I_{0}\setminus\{(p,1),(1,p),(p-1,p-1)\}$. Moreover, the image of $\sigma_{(p,1)}$ and $\sigma_{(1,p)}$ also generates $A^{(1,1)}$. Finally, recall that $\sigma_{(p-1,p-1)}$ coincides with the commutator $\gamma_{1},\gamma_{2}]$ and it generates $A^{(p-1,p-1)}$.
Hence the $\Lambda$-module $A$ is generated by the image of $\{\sigma_{\boldsymbol{m}}\}_{\boldsymbol{m}\in I_{0}}$, which is equivalent to saying that $\operatorname{Gal}(\Omega/K(p^{\infty}))$ is normally generated by $\{\sigma_{\boldsymbol{m}}\}_{\boldsymbol{m}\in I_{0}}$ (as a subgroup of $G$). Hence $\gamma_{1}$, $\gamma_{2}$ and $\{\sigma_{\boldsymbol{m}}\}_{\boldsymbol{m}\in I_{0}\setminus\{(p-1,p-1)\}}$ generate the group $G$, as desired. ∎
@@
Now we recall the discussion occurring at constructing $\sigma_{\boldsymbol{m}}$ and $g_{\boldsymbol{m}}$ for $\boldsymbol{m}\in I$ such that $\boldsymbol{m}\equiv\boldsymbol{1}\bmod p-1$. There were two ways to construct them, starting from $\sigma_{(p,1)}$ and $\sigma_{(1,p)}$, but they yield the same elements on $A=\operatorname{Gal}(\Omega/K(p^{\infty}))^{{\rm ab}}$. Hence, by applying Corollary 4.11 to $\mathcal{F}=G$, $y_{1}=\gamma_{1}$, $y_{2}=\gamma_{2}$ and $\{x_{i}\}_{1\leq i\leq r}=\{g_{\boldsymbol{m}}\}_{\boldsymbol{m}\in I_{0}}$, it follows that $F^{1}G=\operatorname{Gal}(\Omega/K(p^{\infty}))$ is strongly generated by $\{g_{\boldsymbol{m}}\}_{\boldsymbol{m}\in I}$. By Lemma 4.8, the elements $\{\sigma_{\boldsymbol{m}}\}_{\boldsymbol{m}\in I}$ strongly generates $F^{1}G$.
We claim that the surjective map $F^{1}G\to F^{1}G^{\ast}$ is an isomorphism, which is equivalent to the desired equality $\Omega=\Omega^{\ast}$. To prove the claim, let $\mathcal{G}$ be a free pro-$p$ group on the set $\{\tilde{\sigma}_{\boldsymbol{m}}\}_{\boldsymbol{m}\in I}$, $\varphi\colon\mathcal{G}\to F^{1}G$ a surjective homomorphism sending $\tilde{\sigma}_{\boldsymbol{m}}$ to $\sigma_{\boldsymbol{m}}$ for every $\boldsymbol{m}\in I$ and $\varphi^{\ast}\colon\mathcal{G}\to F^{1}G^{\ast}$ the composite of $\varphi$ and $F^{1}G\to F^{1}G^{\ast}$.
The group $\mathcal{G}$ comes equipped with a descending central filtration $\{F^{m}\mathcal{G}\}_{m\geq 1}$ induced by $\{F^{m}G^{\ast}\}_{m\geq 1}$ through the map $\varphi^{\ast}$. Note that the graded Lie algebra associated to $\{F^{m}\mathcal{G}\}_{m\geq 1}$ is the same as $\mathfrak{g}$, and we have
$\operatorname{Ker}(\varphi^{\ast})=(\varphi^{\ast})^{-1}(\bigcap_{m\geq 1}F^{m}G^{\ast})=\bigcap_{m\geq 1}F^{m}\mathcal{G}.$
By Lemma 4.12 (1), the group $F^{m}\mathcal{G}$ contains $\tilde{F}^{m}\mathcal{G}$ for every $m\geq 1$. Hence we have the following commutative diagram:
@@
Since we assume the analogue of the Deligne-Ihara conjecture (Conjecture 2.3), it holds that $\{\tilde{\sigma}_{\boldsymbol{m}}\}_{\boldsymbol{m}\in I}$ freely generates $\bigoplus_{m\geq 1}(F^{m}\mathcal{G}/F^{m+1}\mathcal{G})\otimes\mathbb{Q}_{p}$ as a graded Lie algebra. However, by Lemma 4.12 (2), the Lie algebra $\bigoplus_{m\geq 1}(\tilde{F}^{m}\mathcal{G}/\tilde{F}^{m+1}\mathcal{G})\otimes\mathbb{Q}_{p}$ is also generated by $\{\tilde{\sigma}_{\boldsymbol{m}}\}_{\boldsymbol{m}\in I}$, which forces the upper horizontal arrow in the diagram to be an isomorphism. As a consequence, the lower horizontal arrow in the diagram is found to be injective. As is observed by induction on $m$, this injectivity is equivalent to saying that two filtrations $\{\tilde{F}^{m}\mathcal{G}\}_{m\geq 1}$ and $\{F^{m}\mathcal{G}\}_{m\geq 1}$ coincide with each other. Hence by Lemma 4.12 (3), we have
$\bigcap_{m\geq 1}F^{m}\mathcal{G}=\bigcap_{m\geq 1}\tilde{F}^{m}\mathcal{G}=\{1\},$
showing the injectivity of $\varphi^{\ast}$. Hence the map $\varphi^{\ast}$ is an isomorphism, so is the map $F^{1}G\to F^{1}G^{\ast}$ as desired. ∎
Lastly, we record the following corollary:
Corollary 4.14.
The following assertions hold.
  1. (1)
@@
    The elements $\{\sigma_{\boldsymbol{m}}\}_{\boldsymbol{m}\in I}$ freely generate the integral Lie algebra $\mathfrak{g}$. In other words, under the assumption of Theorem 2.14, the integral version of Conjecture 2.3 holds.
  2. (2)
@@
    The Galois group $\operatorname{Gal}(\Omega/K(p^{\infty}))$ is free pro-$p$ group with $\{\sigma_{\boldsymbol{m}}\}_{\boldsymbol{m}\in I}$ as its basis (note that this result is consistent with Proposition 4.4 (2)).
Proof.
Both assertions follow from the proof of Theorem 2.14 above, together with Lemma 4.12. ∎
Remark 4.14.1.
The integral version of the Deligne-Ihara conjecture is related to generalized Greenberg’s conjecture for $\mathbb{Q}(\mu_{p})$ [Sha02, Theorem 1.3]. Roughly speaking, if the integral version of the Deligne-Ihara conjecture holds, then the image of the pro-$p$ outer Galois representation from $G_{\mathbb{Q}(\mu_{p})}$ is found to be a free pro-$p$ group of rank $\frac{p+1}{2}$. However, such a free pro-$p$ extension cannot exist if $p>2$ is irregular and generalized Greenberg’s conjecture [Gre01, Conjecture (3.5)] holds, by a result of McCallum [McC01, Theorem 2]. Similarly, one can show that the integral version of Conjecture 2.3 also does not hold under similar assumptions. We plan to discuss this topic in a future paper.
Appendix A Pure locality of $p$-th cyclotomic extension of imaginary quadratic field